Status Report

NASA Space Station On-Orbit Status 8 May 2004

By SpaceRef Editor
May 8, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 8 May 2004

All ISS systems continue to function nominally, except those noted previously or below.  Saturday — off duty for the new station crew.

After wakeup and morning inspection, CDR Gennady Padalka, working off the Russian task list, unstowed and set up the equipment for the Cardiocog experiment, then performed the procedure.   [Originally part of Pedro Duque’s VC5 “Cervantes” science program, Cardiocog studies changes in the human cardiovascular system in micro-G, expressed in the peripheral arteries, and the vegetative regulation of arterial blood pressure and heart rate.  For the experiment, the CDR had to take systolic & diastolic blood pressure measurements and pulse data, using the Tensoplus sphygmomanometer, storing the data on the EGE-2 laptop.]

Both crewmembers performed the regular weekly 3-hr. station cleaning, their second.   [“Uborka”, done every Saturday, includes removal of food waste products, cleaning of compartments with vacuum cleaner, wet cleaning of the Service Module (SM) dining table and other surfaces with “Fungistat” disinfectant and cleaning fan screens to avoid temperature rises.]

FE/SO Michael Fincke relocated and set up the radiation measuring instrument TEPC (tissue equivalent proportional counter) at a new station in the Lab near the IVCPDS (intravehicular charged particle directional spectrometer), where it can be checked out against the CPDS data.

Padalka completed the daily routine maintenance of the SM’s SOZh life support system.

After readying the Metox (metal oxide) “bake-out oven” in the Airlock (A/L), Mike Fincke initiated the regeneration process on the last two expended Metox  CO2 adsorbent canisters (#19 & #21), for use in the EMU EVA on 6/10.

During yesterday’s EVA NiMH (Nickel Metal Hydride) batteries discharging activity in the A/L, battery charger #2 (BC2) shut down when its controlling RPC (remote power controller, RPC4), tripped open, interrupting the battery restoration process.  Specialists are discussing possible workarounds and any impacts to the timeline.

Four of the deeply discharged NiMH batteries for the PGT (pistol grip tool) have not accepted a charge, despite four attempts at charging.   [If it turns out that they will not be available for EVA use, the R&R (removal & replacement) of the failed RPCM can not be done with PGTs (the certified torque tool to perform bolt torquing associated with EVA tasks including ISS critical contingencies).  However, a ratchet wrench is available, which could be used for the RPCM R&R task, but before such a workaround is chosen, additional battery recovery steps will probably be attempted.]

At 8:15am EDT, the crew conducted the weekly planning conference with the ground, discussing next week’s “Look-Ahead Plan” (regularly prepared jointly by MCC-H and MCC-M planners), via S-band/audio.

Later, the crew also held their monthly teleconference with ISS Program Management via S-band/audio.

The inspection of the three emergency lighting power sources (ELPS) in the Node scheduled in part for today was all completed yesterday.

Systems monitoring of the Soyuz TMA-4 crew return vehicle (CRV) continues.  Twice today, the CDR again took routine spacecraft parameter readings (Form 03) in the vehicle for calldown to TsUP/Moscow.

Mike Fincke conducted the daily leak check of the Lab window’s inter-pane space (“Volume D”), using the “Aeolus” scopemeter with pressure probe.   [Readings dating back to March confirm a steady leak rate of ~27 Torr (0.52 psi) per day from the cabin into Volume D.  The current pressure checks will continue until 5/21, after which Fincke is to perform leak checks on the pressure test equipment itself, followed by a detailed ULD (ultrasound leak detector) window inspection (to determine if the tiny leaks found previously with the ULD are venting to Volume D or possibly directly to space).  Next will be a fit check of the new protective box coming up on Progress 14P over window ports C & D, and finally installation of the new U-jumper flexhose and protective shield, plus any procedure developed meanwhile for fixing the pinhole leaks.]

Working off the Russian task list, CDR Padalka inspected and set up the equipment for “Diatomeya” ocean observations scheduled for tomorrow and Monday (9/10).   [The gear to be installed at Service Module (SM) windows #7 or #8 include the DVCAM-150 video camera on an articulated bracket and the Nikon F5 electronic still camera with 24/85-mm lens.]

The crew completed their daily 2.5-hr. physical exercise program on TVIS, RED exerciser and VELO cycle with load trainer.   [Engineers continue to investigate the crew-reported problems with the TVIS control panel (resets to zero in motorized mode).  While diagnostics is underway, Mike and Gennady are using the treadmill in passive mode, but with VIS (vibration isolation system) active.]

At 1:00pm EDT, Gennady downlinked a TV message of greeting to Ikram Adirbekovich Adirbekov in Kisil-Ordinsjkoi Province.

Weekly Science Update (Expedition Nine — 2nd):

GASMAP:   “Thank you” to Mike for conducting a successful GASMAP routine health check last week, which helps to maintain the integrity and stability of the machine.

Human Research Facility/Workstation (HRF WS):   Continuing.

Advanced Ultrasound (ADUM):   Mike was thanked for an “excellent job” installing the audio drivers and creatively working around the elusive PuFF data kit for the ADUM OPE installation.   It is hoped that he will “enjoy the OPE session next week”.

Hand Posture Analyzer (HPA:  Nothing new.

In-Space Soldering Investigation (ISSE):   Nothing new.

Pore Formation and Mobility Investigation (PFMI):   Nothing new.
Space Acceleration Measurement System (SAMS):   SAMS telemetry from all sensors is currently not available due to ICU lockup on 5/6 (7:09am EDT).  The ground continues to troubleshoot the ICU recovery.  In the mean time MAMS’s HiRAP is being used to characterize the vibratory environment in the US Lab.
Microgravity Acceleration Measurement System (MAMS):   MAMS telemetry downlink is back up after yesterday’s Express Rack 1 RIC problem.  The MAMS team would like to thank Mike Fincke for his support.  MAMS OSS/HIRAP continue to monitor the microgravity environment.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):  Nominal.

Protein Crystal Growth Monitoring by Digital Holographic Microscope (PromISS):   Nothing new.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE):    Planned.

Binary Colloidal Alloy Test-3 (BCAT-3):   On 5/5, Mike Fincke completed another photo session.  The BCAT-3 team appreciated his willingness for feedback.

Renal Stone (RS):   Nothing new.

Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES):   Nothing new.

Foot/Ground Reaction Forces during Space Flight (FOOT):   Nothing new.

Materials ISS Experiment (MISSE):  In progress. Deployed outside on the U.S. Airlock. Nominal and collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI):   Nothing new.

Yeast Group Activation Packs (Yeast GAP):   Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM):  A new session will start next week, from a Service Module window.

Earth Science Toward Exploration Research (ESTER):  Nothing new.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2):  Planned.

Commercial Generic Bioprocessing Apparatus (CGBA):   Nothing new.

Miscible Fluids in Microgravity (MFMG):    Four runs are remaining; they will be performed on Increment 9.

Educational Payload Operations (EPO):   Nnothing new.

Crew Earth Observations (CEO):  CEO photos are shot with shutter speed set at 1/500sec or higher to stop earth motion, unless lighting conditions dictate a slower shutter speed.  The crew uses manual focus,toobtainthe best possiblefocus, and the NASA-provided 400mm lens forbetter quality telephoto images.
Today’s CEO targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Congo Basin (Dynamic event. Unusually clear weather persists.  Any clearings in the forest are of great interest as logging begins along the fringes of the rainforest, especially right of track in northernmost Rep. Congo [Congo-Brazzaville].  This pass took ISS over the Congo-Ubangi river confluence again.  A detailed shot of the differing water color would have general interest), Savanna fires, Angola, Zambia (Dynamic event.  Thousands of fires are reported as having broken out recently in what is known as the most fire-prone part of the planet.  Thick savanna forests immediately south of the rainforest burn in the dry season.  Biomass in these forests has increased in recent years due to very good rains.  The equatorial concentrations of thunderstorms has made the seasonal shift way north into the southern Sahel [Nigeria].  Plumes are blowing NW.  Oblique views capture the extent of the phenomenon, and any generalized smoke palls as well), and Dakar, Senegal (nadir pass over this city on the spit of land that juts way out into the Atlantic).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

ISS Orbit  (as of this noon, 1:26pm EDT [= epoch]):

  • Mean altitude  360.9 km
  • Apogee — 368.4 km
  • Perigee — 353.5 km
  • Period — 91.76 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0011042
  • Orbits per 24-hr. day — 15.69
  • Mean altitude loss last 24 hours — <50 m
  • Revolutions since FGB/Zarya launch (Nov. 98)  — 31212

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.